OpenStudy (anonymous):

13r + 10 = 5r - 46 r=

OpenStudy (abmon98):

13r-5r=-46-10 8r=-56 r=-56/8

OpenStudy (igreen):

\(13r + 10 = 5r - 46\) \(8r = -56\) \(-7\)

OpenStudy (muzzack):

r = -7

OpenStudy (muzzack):

Step by step solution : Step 1 : Pulling out like terms : 1.1 Pull out like factors : 8r + 56 = 8 • (r + 7) Equation at the end of step 1 : 8 • (r + 7) = 0 Step 2 : Solve 8•(r+7) = 0 Equations which are never true : 2.1 Solve : 8 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : r+7 = 0 Subtract 7 from both sides of the equation : r = -7 One solution was found : r = -7